Solve the following system of equations:
2x + 4y + 6z = 2
x + 2z = 0
2x + 3y – z = -5

Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLE AND TRY.IF STILL IN DIFFICULTY PLEASE COME BACK

Answer 27519 by venugopalramana(2066)on 2006-06-16 10:35:39 
I need help with the addition method. I have a problem that i cannot figure out. Please help.
6x-5y+z=2.............................I
3x-4y+z=2............................II
4x+2y-z=-5...............................III
EQN.I-EQN.II
3X-Y=0........................................IV
EQN.II+EQN.III
7X-2Y=-3.....................................V
2*EQN.IV-EQN.V
-X=3...X=-3
Y=3X=-9
6*-3-5*(-9)+Z=2
Z=-25